Amy Dyken, controller at Oriole Pharmaceutical Industries, a public company, is currently preparing the calculation for basic and diluted earnings per share and the related disclosure for Oriole’s financial statements. Below is selected financial information for the fiscal year ended June 30, 2020.

Oriole Pharmaceutical Industries
Selected Balance Sheet Information
June 30, 2020

Long-term debt
   Notes payable, 9%

$980,000

   8% convertible bonds payable

4,960,000

   9% bonds payable

5,940,000

     Total long-term debt

$11,880,000

Shareholders’ equity
   Preferred stock, 6% cumulative, $50 par value, 105,000 shares authorized, 26,250 shares issued and outstanding

$1,312,500

   Common stock, $1 par, 9,800,000 shares authorized, 980,000 shares issued and outstanding

980,000

   Additional paid-in capital

4,020,000

   Retained earnings

5,930,000

     Total shareholders’ equity

$12,242,500


The following transactions have also occurred at Oriole.

1. Options were granted on July 1, 2019, to purchase 220,000 shares at $16 per share. Although no options were exercised during fiscal year 2020, the average price per common share during fiscal year 2020 was $20 per share.
2. Each bond was issued at face value. The 8% convertible bonds will convert into common stock at 50 shares per $1,000 bond. The bonds are exercisable after 5 years and were issued in fiscal year 2019.
3. The preferred stock was issued in 2019.
4. There are no preferred dividends in arrears; however, preferred dividends were not declared in fiscal year 2020.
5. The 980,000 shares of common stock were outstanding for the entire 2020 fiscal year.
6. Net income for fiscal year 2020 was $1,510,000, and the average income tax rate is 20%.


For the fiscal year ended June 30, 2020, calculate the following for Oriole Pharmaceutical Industries. (Round answers to 2 decimal places, e.g. $2.45.)

(a) Basic earnings per share____

(b) Diluted earnings per share____
Answer #1
Particulars Calculations for the amounts
Calculation of Basic Earning per Share:
Net Income after tax          1,510,000
Less: Preference Dividend              (78,750) = 1312500*.06
Net Income for Common Shares          1,431,250
No. of Common Shares              980,000 = 980000/1
Basic Earnings per share                  1.460
Calculation of Diluted Earning per Share:
Calculation of No. Diluted Earnings:
Net Income for Common Shares          1,510,000 From Part 1 Above
Add: Income of Options (#1)                         -   Nil
Add: Income of Convertible Debentures              317,440 = 4960000 * 0.08 * (1-0.20)
         1,827,440
Calculation of No. Diluted Shares:
Common Shares              980,000 From Part 1 Above
Add: No. of Options (#1)                44,000 = 220000- (220000 * 16 / 20)
Add: No. of Convertible Debentures              248,000 =4960000/1000*50
         1,272,000
Diluted Earning per Shares                  1.437

#1: For calculation of EPS, we calculate the number of options that are issued without any connsideration. Check the calculation as mentioned above. No consideration will be considered for the options so issued.

Final Note:

Since all the securitites are outstanding for the whole year, there is no need to calculate the weighted average number of outstanding securities.